# Deep Learning Foundation and Pattern Models: Challenges in Hydrological Time Series

### v2

@@ 3.1 Long Short-Term Memory
$f_{t}=sigmoid(W_{f}\times x_{t}+U_{f}\times h_{t-1}+b_{f})\vspace{-2mm}$ (3)
$i_{t}=sigmoid(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-2mm}$ (4)
$\hat{C}_{t}=tanh(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-2mm}$ (5)
$O_{t}=sigmoid(W_{o}\times x_{t}+U_{o}\times h_{t-1}+b_{o})\vspace{-2mm}$ (6)
$C_{t}=C_{t-1}\times f_{t}+i_{t}\times\hat{C}_{t}\vspace{-2mm}$ (7)
@@
3.2 Model Setup
The LSTM model used in this study is implemented using the Tensorflow framework, which allows for customization of layer parameters. The baseline model architecture is shown in Fig. 3. Detailed activation function setup is shown in Tab. 4. To prevent overfitting, a dropout rate of 20% is applied to the layers.
@@ 3.2 Model Setup
Table 4: Model Activation Setup
@@
3.3 Model Training and Evaluation
Inputs to the model can be classified into known inputs and observed inputs Fox et al. (2022). Known inputs refer to features that are known in both the past and future. In this hydrology study, static known inputs are the exogenous features such as climatic signatures, hydrologic signatures, and catchment topography. Additionally, time series known inputs exist in the form of seasonal patterns hidden in all time series features. Observed inputs are features known only for the past time periods but unknown in the future. In hydrology, those features include precipitation, temperature, and streamflow. Model output, known as targets, are the time-dependent predicted properties. In this study, predicted targets are precipitation, temperature, and streamflow.
The input data is divided into batches with a sequence length of 21 days, selected after testing various other lengths, including 7, 14, and 365 days. The 3-week sequence length was chosen to effectively capture subtle hydrological patterns. The total number of batches processed during one epoch is calculated using Equation 9, while the batch size is determined by Equation 10. In these equations, $Day_{total}$ represents the total duration of the input data in days, $l_{seq}$ denotes the sequence length, #$\#Gauges$ refers to the total number of gauges in the input data, and #$\#Input\ properties$ indicates the total number of input features, including both static and time series properties.
#$\#batch/epoch=Day_{total}-l_{seq}+1\vspace{-6mm}$ (9)
@@ 3.3 Model Training and Evaluation
Symbolic Window. A key challenge encountered during LSTM model training was the space complexity associated with handling the training data. Under the previously mentioned batch size configuration, storing all batches prior to training requires RAM space of $O(S_{batch}\times\#batch/epoch)$. To address this issue, we leverage a symbolic window to dynamically generate batches during each epoch of training. Instead of storing the entire training set, including all batches, before training, we only store the original input data with a space complexity of $O(Day_{total}\times\#Gauges\times\#Input\ properties)$. During each epoch, we track the start index $i$ for the current batch and extract the corresponding data for the batch from the input data using $Data[day_{i}:day_{i+l_{seq}}]$. This batch is then trained, and the index $i$ is incremented before proceeding to extract the next batch. This process is repeated #$\#batch/epoch$ times to complete a full epoch of training. Note: This technique is environment-dependent and may not be applicable to all frameworks.
Spatial and Temporal Encodings. Like many fields of science, hydrology time series data exhibit strong seasonal patterns. It can reasonably be assumed that, at a specific gauge location, the precipitation and streamflow in October of one year will be similar to those recorded in October of the previous year. Furthermore, research has identified the approximate water residence times in various reservoirs, such as rivers, lakes, soil, and the atmosphere Pidwirny (2006). To effectively capture these known dependencies within the time series properties, spatial and temporal encodings are incorporated into the model during training.
@@
Spatial Validation. We employ location-based rather than temporal-based validation as the catchments in the CAMELS and Caravan datasets are uncorrelated. The training and validation datasets are randomly selected on an 8:2 ratio by location, given the extensive number of catchments included in the datasets used in this study. For example, out of the 671 catchments included in CAMELS-US dataset, 537 catchments are used for training and 134 catchments are used for validation.
NNSE. Model fit is quantified using normalized Nash-Sutcliffe Efficiency (NNSE) scores Nossent and Bauwens (2012), calculated using Equation 11, where $T$ represents the total number of days, $Q_{m}^{t}$ represents the modeled discharge on day $t$, $Q_{o}^{t}$ represents the observed discharge on day $t$, and $\overline{Q}_{o}$ is the mean observed discharge over $T$ days. NNSE values range from 0 (poor performance) to 1 (perfect fit), with a score of 0.5 indicating that the model’s predictions are equivalent to the time-averaged mean of the observations. In this study, the NNSE value is calculated for each gauge over time $T$, and the average NNSE across all gauges in the input data is reported.
@@
4 LSTM Benchmark Runs
4.1 Experiment Setup
All experiments are conducted using the model architecture shown in Fig. 3 and the activation functions listed in Table 4. The benchmark runs utilize Linear Space, Linear Time, and Annual Fourier Time encodings. We leverage multivariate time series forecasting RAMAN and SUNILKUMAR (1995). Streamflow time series data is not trained but predicted as target.
4.2 Three-Nation Combined Benchmark Run
This run evaluates model performance on the combined CAMELS dataset from three nations: the US, UK, and Chile. The input data include static properties that are common across all three regions, whereas predicted targets consist of precipitation, mean temperature, and streamflow. Results are demonstrated in Table 5.
Table 5: CAMELS Three-Nations Combined Benchmark Run
MSE NNSE
@@ 4.2 Three-Nation Combined Benchmark Run
4.3 CAMELS Caravan US Benchmark Runs
The runs compare model performance between the original CAMELS-US dataset and the US sub-dataset within the Caravan dataset. The CAMELS-US model is trained using selected static and time series data from the original CAMELS-US dataset, while the Caravan-US model is trained using selected static and time series data from the US sub-dataset within Caravan. For both runs, we utilize the same time series features—precipitation and mean temperature—for training and predict the same targets: precipitation, mean temperature, and streamflow. Results are presented in Table 6.
Table 6: CAMELS Caravan US Comparison
@@ 4.3 CAMELS Caravan US Benchmark Runs
MSE NNSE MSE NNSE
Train 0.003508 0.820 0.002920 0.851
Precipitation Val 0.003585 0.819 0.004307 0.801
Train 0.000276 0.961 0.000468 0.967
Mean Temperature Val 0.000283 0.960 0.000573 0.963
Train 0.000287 0.806 0.000525 0.814
Streamflow Val 0.000296 0.812 0.000955 0.703
Train 0.004111 - 0.003982 -
Total Val 0.004203 - 0.005987 -
4.4 Caravan PCA Runs
This experiment explores the use of Principal Component Analysis (PCA) Wold et al. (1987) to reduce the dimensionality of static properties in the Caravan input data. The Caravan sub-datasets contain over 200 static properties, nearly seven times more than the number of static input properties used in the original CAMELS studies. While this extensive range of static properties enhances model training, it significantly increases computational demands and GPU usage. To address this, we apply PCA, a widely adopted dimensionality reduction technique, to reduce the number of static input features to a level comparable with that in the CAMELS studies. In this experiment, we set the explained variance threshold to 90%, resulting in the reduction of static input features to approximately 30.
The first part of the experiment assesses the effect of PCA on model trained on US sub-dataset within the Caravan dataset, representing a small-scale input. The second part of the experiment examines the effect of PCA on model trained on North America regional data (HYSETS) Arsenault et al. (2020) within the Caravan dataset, representing a large-scale input.
Results, shown in Table 7 and Table 8, indicate that the models trained with static properties obtained from PCA perform comparably to the models trained with original static properties, demonstrating that the reduction in input static dimensionality does not significantly compromise model accuracy. These findings validate PCA as an effective approach for train hydrology time series models with high static dimensionality.
Table 7: Caravan US PCA Experiment
@@ 4.4 Caravan PCA Runs
Table 8: Caravan Hysets PCA Experiment
@@
MSE NNSE MSE NNSE
Train 0.002714 0.835 0.002809 0.830
Precipitation Val 0.002960 0.826 0.003055 0.821
Train 0.000372 0.968 0.000386 0.967
Mean Temperature Val 0.000385 0.966 0.000397 0.965
Train 0.000614 0.825 0.000699 0.813
Streamflow Val 0.000876 0.798 0.000397 0.799
Train 0.003470 - 0.003611 -
Total Val 0.003897 - 0.003926 -
4.5 Caravan Global Benchmark Run
This study provides insights into the model’s applicability to large-scale global hydrology datasets. The input data is compiled by concatenating all seven Caravan sub-datasets, which encompass catchments from four continents. As shown in Table 9, accuracy decreases slightly compared to small-scale, individual nation fits (Table 6), yet overall model performance remains relatively high, demonstrating its robustness on global datasets.
Table 9: Caravan Global Fit
MSE NNSE
Train 0.003115 0.809
Precipitation Val 0.003223 0.808
Train 0.000357 0.953
Mean Temperature Val 0.000357 0.952
Train 0.000675 0.781
Streamflow Val 0.000740 0.768
Train 0.003965 -
Total Val 0.004133 -
@@ 4.5 Caravan Global Benchmark Run
5 Static Properties and Spatial Temporal Encodings Experiment
Figure 4: CAMELS US experiment (MSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 5: CAMELS three-nations combined experiment (MSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 6: Caravan US experiment (MSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 7: Caravan combined experiment (MSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
5.1 Experiment Setup
This set of experiments examines the influence of static properties and spatial-temporal encodings on rainfall-runoff modeling accuracy, demonstrated through eight run configurations on four datasets. The first experiment uses the original CAMELS-US dataset Addor et al. (2017), representing a traditional, locally-sourced small-scale dataset. Results are presented in Fig. 7. The second experiment utilizes the Three-Nation Combined CAMELS dataset Addor et al. (2017); Coxon et al. (2020); Alvarez-Garreton et al. (2018), representing a locally-sourced mid-scale dataset. Results are shown in Fig. 7. The third experiment is conducted on the US sub-dataset of the Caravan dataset Kratzert et al. (2023), representing a globally-sourced small-scale dataset. Results are shown in Fig. 7. The fourth experiment uses the full Caravan concatenated dataset Kratzert et al. (2023), representing a globally-sourced large-scale dataset. Results are shown in Fig. 7. All four spatial-temporal encoding configurations tested are presented in Table 10. For each configuration, we conduct runs with and without static features as input to highlight the impact of static features. The model architecture remained consistent across all runs.
Table 10: Static Properties and Encodings Experiment Configurations
@@ 5.2 Experiment Findings
6 Data Splitting and Cube Root Normalization Experiment
Table 11 summarizes four experimental configurations for training and evaluating the LSTM model on precipitation, varying the train/validation set splitting (time-based vs. location-based) and preprocessing (just min-max vs. min-max and cube root normalization). The results, sorted by the Normalized Nash–Sutcliffe Efficiency (NNSE), shed light on how domain-specific considerations can significantly affect model performance.
Table 11: Performance of LSTM across different data preprocessing and splitting strategies.
Splitting Preprocessing MSE MAE NNSE
Time min-max 0.00035 0.00949 0.6177
Time min-max & cube root 0.00682 0.06076 0.7180
Location min-max 0.00020 0.00707 0.7443
Location min-max & cube root 0.00365 0.04121 0.8240
Location-based splitting with cube root normalization achieves the highest NNSE (0.8240), demonstrating that LSTMs perform better in modeling rainfall-runoff relationships for unseen catchments when precipitation data are cube-root-transformed. This transformation reduces skewness caused by large variations in precipitation across catchments with differing climates and land surfaces, enabling better learning from extreme values. In comparison, using raw data with the same location-based split results in a lower NNSE (0.7443), likely due to skewed values disrupting stable training across diverse catchments.
@@ 6 Data Splitting and Cube Root Normalization Experiment
MSE and MAE, reported in Table 11, vary across test sets (distinct years for time-based, distinct catchments for location-based) and are not directly comparable. NNSE, however, normalizes residuals by the observed variance, providing a consistent and reliable measure of model performance across different configurations. This makes NNSE the most appropriate metric for evaluating how well the model captures hydrological dynamics.
7 Comparison with other Pattern and Foundation Models

### v3

@@ III-A Long Short-Term Memory
$f_{t}=sigmoid(W_{f}\times x_{t}+U_{f}\times h_{t-1}+b_{f})\vskip-5.69054pt$ (3)
$i_{t}=sigmoid(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vskip-5.69054pt$ (4)
$\hat{C}_{t}=tanh(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vskip-5.69054pt$ (5)
$O_{t}=sigmoid(W_{o}\times x_{t}+U_{o}\times h_{t-1}+b_{o})\vskip-5.69054pt$ (6)
$C_{t}=C_{t-1}\times f_{t}+i_{t}\times\hat{C}_{t}\vskip-5.69054pt$ (7)
@@
III-B Model Setup
The LSTM model used in this study is implemented using the Tensorflow framework, which allows for customization of layer parameters. The baseline model architecture is shown in Fig. 3. Detailed activation function setup is shown in Tab. IV. To prevent overfitting, a dropout rate of 20% is applied to the layers.
@@ III-B Model Setup
TABLE IV: Model Activation Setup
@@
III-C Model Training and Evaluation
Inputs to the model can be classified into known inputs and observed inputs [5]. Known inputs refer to features that are known in both the past and future. In this hydrology study, static known inputs are the exogenous features such as climatic signatures, hydrologic signatures, and catchment topography. Additionally, time series known inputs exist in the form of seasonal patterns hidden in all time series features. Observed inputs are features known only for the past time periods but unknown in the future. In hydrology, those features include precipitation, temperature, and streamflow. Model output, known as targets, are the time-dependent predicted properties. In this study, predicted targets are precipitation, temperature, and streamflow.
The input data is divided into batches with a sequence length of 21 days, selected after testing various other lengths, including 7, 14, and 365 days. The 3-week sequence length was chosen to effectively capture subtle hydrological patterns. The total number of batches processed during one epoch is calculated using Equation 9, while the batch size is determined by Equation 10. In these equations, $Day_{total}$ represents the total duration of the input data in days, $l_{seq}$ denotes the sequence length, #Gauges\#Gauges refers to the total number of gauges in the input data, and #Inputproperties\#Input\ properties indicates the total number of input features, including both static and time series properties.
$batch/epoch=Day_{total}-l_{seq}+1\vskip-17.07164pt$ (9)
@@ III-C Model Training and Evaluation
Symbolic Window. A key challenge encountered during LSTM model training was the space complexity associated with handling the training data. Under the previously mentioned batch size configuration, storing all batches prior to training requires RAM space of $O(S_{batch}\times\#batch/epoch)$. To address this issue, we leverage a symbolic window to dynamically generate batches during each epoch of training. Instead of storing the entire training set, including all batches, before training, we only store the original input data with a space complexity of $O(Day_{total}\times\#Gauges\times\#Input\ properties)$. During each epoch, we track the start index $i$ for the current batch and extract the corresponding data for the batch from the input data using $Data[day_{i}:day_{i+l_{seq}}]$. This batch is then trained, and the index $i$ is incremented before proceeding to extract the next batch. This process is repeated #batch/epoch\#batch/epoch times to complete a full epoch of training. Note: This technique is environment-dependent and may not be applicable to all frameworks.
Spatial and Temporal Encodings. Like many fields of science, hydrology time series data exhibit strong seasonal patterns. It can reasonably be assumed that, at a specific gauge location, the precipitation and streamflow in October of one year will be similar to those recorded in October of the previous year. Furthermore, research has identified the approximate water residence times in various reservoirs, such as rivers, lakes, soil, and the atmosphere [52]. To effectively capture these known dependencies within the time series properties, spatial and temporal encodings are incorporated into the model during training.
@@
Spatial Validation. We employ location-based rather than temporal-based validation as the catchments in the CAMELS and Caravan datasets are uncorrelated. The training and validation datasets are randomly selected on an 8:2 ratio by location. For example, out of the 671 catchments included in CAMELS-US dataset, 537 catchments are used for training and 134 catchments are used for validation. This decision is made given the extensive number of catchments available in the datasets used in this study.
Training Details. Each benchmark model is trained for 120 successful epochs, a value empirically selected to balance predictive accuracy and computational efficiency. An epoch is considered successful if either the training or validation loss improves relative to the previous epoch. We employ the Adam optimizer [53] with a learning rate of 0.001. Final model evaluation is performed using the checkpoint from the 120th epoch if the loss curve indicates no signs of overfitting. All training is performed on an NVIDIA A100 GPU via Google Colab. On average, training a single epoch with the CAMELS-US dataset [27] requires approximately 30 minutes, while an epoch with the Caravan global dataset takes about one hour. With the same setup, GPU memory usage peaks at 1811MB for CAMELS-US run and 9431MB for Caravan global run.
Evaluation Metrics. Model fit is quantified using root mean squared error (RMSE) and normalized Nash-Sutcliffe Efficiency (NNSE) scores [54]. The NNSE score is calculated using Equation 11, where $T$ represents the total number of days, $Q_{m}^{t}$ represents the modeled discharge on day $t$, $Q_{o}^{t}$ represents the observed discharge on day $t$, and $\overline{Q}_{o}$ is the mean observed discharge over $T$ days. NNSE values range from 0 (poor performance) to 1 (perfect fit), with a score of 0.5 indicating that the model’s predictions are equivalent to the time-averaged mean of the observations. In this study, the NNSE value is calculated for each gauge over time $T$, and the average NNSE across all gauges in the input data is reported.
@@
IV LSTM Benchmark Runs
IV-A Experiment Setup
All experiments are conducted using the model architecture shown in Fig. 3 and the activation functions listed in Table IV. The benchmark runs utilize Linear Space, Linear Time, and Annual Fourier Time encodings. We leverage multivariate time series forecasting [55]. Streamflow time series data is not trained but predicted as target.
IV-B Three-Nation Combined Benchmark Run
This run evaluates model performance on the combined CAMELS dataset from three nations: the US, UK, and Chile. The input data include static properties that are common across all three regions, whereas predicted targets consist of precipitation, mean temperature, and streamflow. Results are demonstrated in Table V.
TABLE V: CAMELS Three-Nations Combined Benchmark Run
RMSE NNSE
@@ IV-B Three-Nation Combined Benchmark Run
IV-C CAMELS Caravan US Benchmark Runs
The runs compare model performance between the original CAMELS-US dataset and the US sub-dataset within the Caravan dataset. The CAMELS-US model is trained using selected static and time series data from the original CAMELS-US dataset, while the Caravan-US model is trained using selected static and time series data from the US sub-dataset within Caravan. For both runs, we utilize the same time series features—precipitation and mean temperature—for training and predict the same targets: precipitation, mean temperature, and streamflow. Results are presented in Table VI.
TABLE VI: CAMELS Caravan US Comparison
@@ IV-C CAMELS Caravan US Benchmark Runs
RMSE NNSE RMSE NNSE
Train 0.0592 0.820 0.0540 0.851
Precipitation Val 0.0599 0.819 0.0656 0.801
Train 0.0166 0.961 0.0216 0.967
Mean Temperature Val 0.0168 0.960 0.0239 0.963
Train 0.0169 0.806 0.0229 0.814
Streamflow Val 0.0172 0.812 0.0309 0.703
Train 0.0641 - 0.0631 -
Total Val 0.0648 - 0.0774 -
IV-D Caravan PCA Runs
This experiment explores the use of Principal Component Analysis (PCA) [56] to reduce the dimensionality of static properties in the Caravan input data. The Caravan sub-datasets contain over 200 static properties, nearly seven times more than the number of static input properties used in the original CAMELS studies. While this extensive range of static properties enhances model training, it significantly increases computational demands and GPU usage. To address this, we apply PCA, a widely adopted dimensionality reduction technique, to reduce the number of static input features to a level comparable with that in the CAMELS studies. In this experiment, we set the explained variance threshold to 90%, resulting in the reduction of static input features to approximately 30.
The first part of the experiment assesses the effect of PCA on model trained on US sub-dataset within the Caravan dataset, representing a small-scale input. The second part of the experiment examines the effect of PCA on model trained on North America regional data (HYSETS) [43] within the Caravan dataset, representing a large-scale input.
Results, shown in Table VII and Table VIII, indicate that the models trained with static properties obtained from PCA perform comparably to the models trained with original static properties, demonstrating that the reduction in input static dimensionality does not significantly compromise model accuracy. These findings validate PCA as an effective approach for train hydrology time series models with high static dimensionality.
TABLE VII: Caravan US PCA Experiment
@@ IV-D Caravan PCA Runs
TABLE VIII: Caravan Hysets PCA Experiment
@@
RMSE NNSE RMSE NNSE
Train 0.0521 0.835 0.0530 0.830
Precipitation Val 0.0544 0.826 0.0553 0.821
Train 0.0193 0.968 0.0196 0.967
Mean Temperature Val 0.0196 0.966 0.0199 0.965
Train 0.0248 0.825 0.0264 0.813
Streamflow Val 0.0296 0.798 0.0199 0.799
Train 0.0589 - 0.0601 -
Total Val 0.0624 - 0.0627 -
IV-E Caravan Global Benchmark Run
This study provides insights into the model’s applicability to large-scale global hydrology datasets. The input data is compiled by concatenating all seven Caravan sub-datasets, which encompass catchments from four continents. As shown in Table IX, accuracy decreases slightly compared to small-scale, individual nation fits (Table VI), yet overall model performance remains relatively high, demonstrating its robustness on global datasets.
TABLE IX: Caravan Global Fit
RMSE NNSE
Train 0.0558 0.809
Precipitation Val 0.0568 0.808
Train 0.0189 0.953
Mean Temperature Val 0.0189 0.952
Train 0.0260 0.781
Streamflow Val 0.0272 0.768
Train 0.0630 -
Total Val 0.0643 -
@@ IV-E Caravan Global Benchmark Run
V Static Properties and Spatial Temporal Encodings Experiment
Figure 4: CAMELS US experiment (RMSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 5: CAMELS three-nations combined experiment (RMSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 6: Caravan US experiment (RMSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
Figure 7: Caravan combined experiment (RMSE loss). Each run configuration is utilized in 2 runs (left: static features not trained, right: static features trained).
V-A Experiment Setup
This set of experiments examines the influence of static properties and spatial-temporal encodings on rainfall-runoff modeling accuracy, demonstrated through eight run configurations on four datasets. The first experiment uses the original CAMELS-US dataset [27], representing a traditional, locally-sourced small-scale dataset. Results are presented in Fig. 7. The second experiment utilizes the Three-Nation Combined CAMELS dataset [27, 28, 29], representing a locally-sourced mid-scale dataset. Results are shown in Fig. 7. The third experiment is conducted on the US sub-dataset of the Caravan dataset [42], representing a globally-sourced small-scale dataset. Results are shown in Fig. 7. The fourth experiment uses the full Caravan concatenated dataset [42], representing a globally-sourced large-scale dataset. Results are shown in Fig. 7. All four spatial-temporal encoding configurations tested are presented in Table X. For each configuration, we conduct runs with and without static features as input to highlight the impact of static features. The model architecture remained consistent across all runs.
TABLE X: Static Properties and Encodings Experiment Configurations
@@ V-B Experiment Findings
VI Data Splitting and Cube Root Transformation Experiment
Table XI summarizes four experimental configurations for training and evaluating the LSTM model on precipitation, varying the train/validation set splitting (time-based vs. location-based) and preprocessing (just min-max scaling vs. min-max scaling and cube root transformation). The results, sorted by the Normalized Nash–Sutcliffe Efficiency (NNSE), shed light on how domain-specific considerations can significantly affect model performance. All runs from this experiment use the baseline LSTM from Nixtla’s NeuralForecast framework [57], as the custom LSTM model introduced in this study currently supports only location-based splits.
TABLE XI: Performance of LSTM across different data preprocessing and splitting strategies.
Splitting Preprocessing RMSE NNSE
Time min-max 0.0187 0.618
Time min-max & cube root 0.0826 0.718
Location min-max 0.0141 0.744
Location min-max & cube root 0.0604 0.824
Location-based splitting with cube root transformation achieves the highest NNSE (0.8240), demonstrating that LSTMs perform better in modeling rainfall-runoff relationships for unseen catchments when precipitation data are cube-root-transformed. This transformation reduces skewness caused by large variations in precipitation across catchments with differing climates and land surfaces, enabling better learning from extreme values. In comparison, using raw data with the same location-based split results in a lower NNSE (0.7443), likely due to skewed values disrupting stable training across diverse catchments.
@@ VI Data Splitting and Cube Root Transformation Experiment
RMSE results, reported in Table XI, vary across test sets (distinct years for time-based, distinct catchments for location-based) and are not directly comparable. NNSE, however, normalizes residuals by the observed variance, providing a consistent and reliable measure of model performance across different configurations. This makes NNSE the most appropriate metric for evaluating how well the model captures hydrological dynamics.
TABLE XII: Performance comparison of various models
Model Type a Input&Output b P c SR d Tmax e Tmin f VP g Q h
LSTM-multivariate P MV&MV 0.0546 0.0509 0.0138 0.0197 0.0122 0.0161
LSTM-univariate P UV&UV 0.0583 0.0567 0.0195 0.0195 0.0155 0.0190
TFT P UV&UV 0.0953 0.0918 0.0277 0.0305 0.0257 0.2385
TSMixer-M4 F MV&MV 0.0849 0.0774 0.0288 0.0226 0.0311 0.0648
TSMixer-TrafficL F MV&MV 0.0859 0.0780 0.0293 0.0230 0.0318 0.0210
TSMixer F MV&MV 0.0931 0.0835 0.0311 0.0270 0.0366 0.0255
TCN P UV&UV 0.1036 0.0943 0.0344 0.0329 0.0383 0.0535
iTransformer-M4 F MV&MV 0.0892 0.0863 0.0354 0.0283 0.0356 0.0202
iTransformer F MV&MV 0.0869 0.0881 0.0356 0.0257 0.0351 0.0232
DilatedRNN P UV&UV 0.0993 0.0947 0.0358 0.0313 0.0375 0.0279
iTransformer-TrafficL F MV&MV 0.0920 0.0889 0.0374 0.0307 0.0373 0.0205
GNNCoder-univariate P UV&UV 0.0843 0.0884 0.0389 0.0345 0.0395 0.0200
PatchTST-Weather F UV&UV 0.0944 0.0920 0.0411 0.0335 0.0405 0.0207
PatchTST-TrafficL F UV&UV 0.0943 0.0920 0.0411 0.0333 0.0406 0.0207
PatchTST-M4 F UV&UV 0.0944 0.0920 0.0411 0.0333 0.0407 0.0207
GNNCoder-multivariate P MV&MV 0.0928 0.0407 0.0414 0.0893 0.0497 0.0200
TimesNet P UV&UV 0.0969 0.0934 0.0420 0.0345 0.0421 0.2943
PatchTST F UV&UV 0.0976 0.0932 0.0429 0.0335 0.0409 0.0235
TiDE P UV&UV 0.1014 0.0933 0.0452 0.0400 0.0518 0.0310
Chronos-T5 F UV&UV 0.1895 0.1236 0.0792 0.0753 0.0851 0.1403
Chronos-GPT2 F UV&UV 0.1675 0.1123 0.0794 0.0814 0.1601 0.1193
Table shows root mean squared error results, with the lowest for each predicted target highlighted in red.
VII Comparison with other Pattern and Foundation Models